# What does really matter in image goal navigation?

### v1

@@ Abstract
Image goal navigation requires two different skills: firstly, core navigation skills, including the detection of free space and obstacles, and taking decisions based on an internal representation; and secondly, computing directional information by comparing visual observations to the goal image. Current state-of-the-art methods either rely on dedicated image-matching, or pre-training of computer vision modules on relative pose estimation. In this paper, we study whether this task can be efficiently solved with end-to-end training of full agents with RL, as has been claimed by recent work. A positive answer would have impact beyond Embodied AI and allow training of relative pose estimation from reward for navigation alone. In a large study we investigate the effect of architectural choices like late fusion, channel stacking, space-to-depth projections and cross-attention, and their role in the emergence of relative pose estimators from navigation training. We show that the success of recent methods is influenced up to a certain extent by simulator settings, leading to shortcuts in simulation. However, we also show that these capabilities can be transferred to more realistic setting, up to some extend. We also find evidence for correlations between navigation performance and probed (emerging) relative pose estimation performance, an important sub skill.
@@ 1 Introduction
Early work addressed the problem in a purely data-driven way by end-to-end training Late Fusion (eg. Siamese) image encoders [77], with mixed results. It was conjectured, that the weak and sparse learning signal from RL training could potentially not be sufficient to train high capacity encoders comparing images. Since the ImageNav task introduces a sub task of relative pose estimation (“where is the goal with respect to my current position?”), there is a strong case for modular methods separating perception and decision taking. For instance, Krantz et al. [38] employ local features and matching (SuperGlue [58]) combined with map and plan strategies for navigation. However, it has since been shown [8] that image correspondences can actually emerge directly in the cross-attention layers of a binocular transformer, provided it is appropriately pre-trained. This strategy, used in the DEBiT agent [8] seems to outperform the explicit matching approaches. As reported in [8], key to this solution is, firstly, pre-training, and secondly, the fusion of observation and goal in such a way that comparisons between local parts (e.g. patches) of the two respective images can be calculated by the architecture, as opposed to comparing embedding vectors.
More recently, some work seems to indicate that ImageNav can be addressed by very low-capacity convolutional encoders (ResNet9) with channel-wise stacking and through RL-training alone, without any pre-training. The FGPrompt model [67] achieves results which are almost on par with DEBiT with a fraction of the model capacity, which raises interesting questions: can directional information be learned by RL losses alone, and by comparably simple architectures? The implications could potentially go beyond image goal navigation and have impact on the wider field of relative pose estimation and localization.
In this paper we disentangle the different mediators and ask fundamental questions on what choices matter in this topic. We find that (i) Certain network architectures better support implicit correspondence computations, that this directly impacts results, with early patch-wise fusion being essential, compared to late fusion. (ii) We show that recent success of frugal architectures using channel stacking is linked to a simulator setting that allows agents to slide along walls. Interestingly, while this setting is known to hamper real-world transfer, to our surprise, some learned capabilities can still be transferred to realistic settings if the transfer includes weights of the perception module. (iii) We also find correlations between navigation and (emergent) relative pose estimation performance, which we probe from representations trained with RL.
@@ 2 Related Work
Navigation from visual observations – has been addressed in multiple different fields. Robotics for a long time focused on explicit modeling [13, 43, 45], which is mainly based on mapping and localization [12, 39, 68], explicit planning [37, 62] and low-level control [26, 57]. Accurate sensor and observation models are essential, followed by filtering, dynamical models and optimization techniques.
Modern AI-based solutions are typically trained on large-scale photo-realistic simulators like Habitat [59] or AI2-Thor [36]. Modular agents [17] decompose the problem in sub modules, typically mapping, pose-estimation, planning and local decision taking. On the other end of this spectrum, end-to-end trained models directly map input to actions, with Reinforcement Learning (RL) [31, 48, 76, 69] or Imitation Learning (IL) [23], offline-RL [66], or unsupervised RL [35, 42]. In recent work, end-to-end trained agents are combined with visual foundation models, either general ones like DINOv2 [50] in PoliFormer [76], or binocular ones for image comparisons as in DEBiT [8]. Agent memory is typically implemented as flat recurrent representations [10, 32], occupancy maps [17], semantic maps [16], latent metric maps [5, 30, 51], topological maps [6, 18, 63], scene graphs [65], explicit episodic memory [20, 24, 25, 54], implicit representations [46] or navigability [9]. In this work we investigate agents trained end-to-end with RL and recurrent memory.
@@
Relative pose estimation – was tackled for decades with pixel level image matching techniques [28, 49, 60]. Learning-based approaches have also been proposed [34]: to overcome the lack of generalizability of the early methods, self-supervised techniques have been used [41]. More recently, DUSt3R [70] regresses pointmaps of each image expressed in the coordinate system of the first image while MASt3R [40] additionally learns an extra descriptor inspired by standard image matching. Both leverage CroCo as pre-training [71, 72]. Similarly, MicKey [4] regresses pointmaps of each image in its respective coordinate system and learns descriptors to obtain correspondences: this is supervised by relative pose alone with differentiable RANSAC formulations [7, 11]. All these recent methods have led to impressive results for relative pose estimation, even under scenarios with little overlap between input images as in the MapFree-Relocalization benchmark [3]. In this work, we study rel. pose estimation as a sub-task of ImageNav and probe it explicitly from agent representations.
@@ 3 Can we learn goal-perception with actions?
In this work we study the ImageNav task in photo-realistic 3D environments, where an agent is given a goal image $\mathbf{g}{\in}\mathbb{R}^{3{\times}H{\times}W}$ and is required to navigate from a starting location to the position shown in the goal. At each timestep $t$ the agent observes an RGB image $\mathbf{o}_{t}{\in}\mathbb{R}^{3{\times}H{\times}W}$, all images are of size $112{\times}112$. The action space is the discrete set $\mathcal{A}$ ={MOVE FORWARD 0.25m, TURN LEFT $10^{\circ}$, TURN RIGHT $10^{\circ}$, and STOP}. An episode is considered successful if the agent is within 1m of the goal position and it calls the STOP action within its 1000 steps budget. We use the Habitat simulator [59] and the Gibson dataset [74].
@@
$\begin{array}[]{lll}\tilde{\mathbf{g}}_{t}&=\phi(\mathbf{o}_{t},\mathbf{g})&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// binocular perception}}\\ \mathbf{h}_{t}&=h(\mathbf{h}_{t-1},\tilde{\mathbf{g}}_{t},\zeta(\mathbf{a}_{t-1}))&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// state update}}\\ \mathbf{a}_{t}&\sim\pi(\mathbf{h}_{t}),&\texttt{\scriptsize{\color[rgb]{0.96875,0.7734375,0.65234375}\definecolor[named]{pgfstrokecolor}{rgb}{0.96875,0.7734375,0.65234375}// policy}}\\ \end{array}$ (1)
where $h$ is the function updating the hidden state $\mathbf{h}_{t}$ of a GRU [21] and gating equations have been omitted for brevity; $\phi$ and $\zeta$ are trainable encoders, and $\pi$ is a linear policy. All agent variants are trained from scratch with PPO [61] with a reward function defined as in [71], which re-uses the definition from [19] for PointGoal,
$r_{t}=\mathrm{K}\cdot\mathbf{1}_{\text{success}}-\Delta_{t}^{\mathrm{Geo}}-\lambda,$
where $\lambda{=}0.01$ is a slack cost to encourage efficiency, $K{=}10$, and $\Delta_{t}^{\mathrm{Geo}}$ is the increase in geodesic distance to the goal.
@@ 3.1 Architecture design choices
At the heart of the problem is the sub task of learning to compare images $\mathbf{o}_{t}$ and $\mathbf{g}$ to infer directional information related to relative pose estimation. In the classical ImageNav task, goal images can be taken from any position and do not necessarily show a semantically meaningful object. Given the possible absence of semantic cues, which would have allowed to compare images on a global level, we conjecture that successful comparison requires the computation of some form of correspondence between local parts of the respective images. While pose estimation methods based on local feature matching perform this explicitly [55, 58], recent work have shown that this kind of correspondence can emerge implicitly from large-scale training without explicit supervision of matches, e.g. in CroCo [71] for pairwise problems or DUSt3R [70] involving multiple images and addressing SfM tasks.
@@
Starting from the premise, that correspondence computations are necessary to extract directional information, we conjecture that these computations need to be supported by the underlying network architectures. We will first describe the basic building blocks for image comparison presented in the literature, and also illustrated in Figure 2:
@@
use separate networks $\phi_{o}$ and $\phi_{g}$ to encode observation and goal. The representation fed to the agent is given as $\phi(\mathbf{o}_{t},\mathbf{g})=[\phi_{o}(\mathbf{o}_{t}),\phi_{g}(\mathbf{g})]$, where [.][.][ . ] denotes concatenation, see Figure 2(a). The comparison is thus performed “late” between embeddings $\phi_{o}(\mathbf{o}_{t})$ and $\phi_{g}(\mathbf{g})$, which makes it generally harder to be done on a local image level, unless output representations retain sufficient spatial structure going through encoder layers.
@@
involves the usage of a single network to encode both, observation and goal, which are channel stacked into one input image, denoted as $\phi\left([\mathbf{o}_{t},\mathbf{g}]_{\textrm{dim}=1}\right)$, where 0 is the batch dimension and 1 the channel dimension — see Figure 2(b).
If encoders are implemented as CNNs, a common choice in robotics for computational complexity reasons, comparisons between local parts of the respective images can be done only if the receptive field of the network is big enough to encompass the full image. Given a sufficiently deep encoder, correspondence can therefore be computed, but not necessarily efficiently, as individual matching results need to be accumulated in the output of different filter kernels.
If encoders are ViTs, each feature patch corresponds to a pair of patches having the same coordinates in both images. Correspondence can be more easily computed if different self-attention heads are learned with query and key projections focusing on one of the two channel groups of the two respective images [^2] As an illustrative example, consider an input image $X$ corresponding to channel stacking of two images $X_{A}$ and $X_{B}$. Then, if for a given attention head the Query projection only uses the channels of image $X_{A}$ and the Key projection only use the channels of image $X_{B}$, then this self-attention on $X$ mimics cross-attention between $X_{A}$ and $X_{B}$..
@@
reshapes image patches into channel values and was introduced in  [56]. FGPrompt [67] uses SpaceToDepth in combination with Channel stacking observation and goal, achieving strong navigation results. We will investigate whether a ResNet in this configuration can compute correspondences across a large spatial dimension through a single convolutional layer, somewhat reminiscent of cross-attention, only with very few parameters — see Figure 2(c).
@@
is a natural way to compute correspondences between local parts of images [8], as each patch in one image can be naturally linked to one or more patches in the other image through the cross-attention distribution, see Figure 2(d).
@@ 3.2 Implementations in prior work
Early works used late fusion approaches based on CNNs [1, 77], but later work switched encoders to ViTs, for instance OVRL-v2 [75]. Cross-attention for ImageNav was introduced in the DEBiT encoder [8], with code made available [^3] https://github.com/naver/debit, and combined with pre-training, first for cross-view completion as in [71], then for relative pose and visibility estimation. The binocular ViT was combined with an additional CNN taking only the observation as input, trained with RL. It currently provides the state-of-the-art results on this task. Interestingly, the authors have shown that pre-training is essential, as the high-capacity binocular transformer cannot be trained from scratch with RL alone, leading to performance close to zero. The RL loss seems to be too weak of a learning signal, not sufficient to drive the encoder to discover correspondence computations. The inability of RL to learn big encoders was somewhat confirmed in the Late Fusion method OVRL-v2 [75], as “switching from the 50.9M parameters ViT-Small to the 179.2M parameters ViT-Base produces another negative result: SR drops -2.4% while SPL minimally increases by +0.7%”.
FGPrompt [67] first introduced channel stacking for ImageNav, with code made available [^4] https://github.com/XinyuSun/FGPrompt. It is heavily based on the previous code of ZSEL [1]. Both codebases use a non-standard ResNet9 architecture that adopts several custom blocks, the most significant one being the SpaceToDepth modification, described earlier. Interestingly, SpaceToDepth was first introduced in TResNet [56] as a general purpose network block for computer vision optimized for memory efficiency, targeting image classification and object detection. Correspondence problems were not explicitly targeted. The ResNet9 encoder in this paper is exactly the same as used in ZSEL and FGPrompt.
Given the good results FGPrompt obtained without any pre-training, 92% SR, vs. DEBiT’s 94% SR depending on pre-training, we raise the question whether it is possible to use the weak learning signal provided by RL to learn binocular perception modules capable of providing directional information. A positive answer would have significant impact also on training vision modules targeting RPE directly. In the experimental section we will dive into this question and benchmark architectures and design choices.
@@ 4 Experimental Setup
Unfortunately there are multiple different evaluation protocols in the literature on Embodied AI, and in particular on the ImageNav task. We therefore train all agents in identical conditions and with the same experimental protocol. The agents are trained on the 72 scenes of the Gibson dataset [74] and use the standard Habitat episode definitions (a subset of the literature uses the definitions from Episodic Transformers [47] and another subset uses the definitions from NSNRL [27]). We set the maximum episode length to the default value of 1000 steps and we train for 500M steps, which allows all agent variants to converge easily.
@@
Implementation details – We test perception networks $\phi$ implemented as ResNet, ViT [^5] https://github.com/huggingface/pytorch-image-models and DEBiT architectures, available in their respective public repositories. For the ResNet, we consider a ResNet9 because of strong performance reported in [67], close to those obtained with a ResNet50, but with a fraction of the parameters. We use the Small version of ViT since [75] reports higher navigation performance than Base (when not pre-trained, as here), again with a fraction of the parameters. The function $\zeta$ encodes previous actions into a 32D feature, $h$ is a GRU with 2 layers and hidden dimension 128, followed by a linear Actor-Critic policy $\pi$ — see the appendix for more details.
@@
Validation – again, there does not seem to be a clear protocol for choosing checkpoints from a training run, as some papers seem to take the last and some report the best validation performance over several checkpoints, sometimes up to several thousands. To keep the experimental protocol sound and avoid overfitting evaluation over the choice of checkpoints, while at the same time staying comparable, we use three different splits. We train on the standard Gibson train split (72 scenes) and report on the standard Gibson validation split (14 scenes), as the literature, but we use an independent custom holdout set of scenes to choose the checkpoint. These additional scenes have been taken from the unused set of the full Gibson set.
Model s2d Backbone PT SR(%) SPL(%)
(a) Late Fusion ResNet9 13.8 8.0
(b) [1] Late Fusion ResNet9 12.5 7.6
(c) Late Fusion ViT-Small 12.5 6.7
(d) ChannelCat ResNet9 83.2 43.9
(e) [67] ChannelCat ResNet9 83.6 42.1
(f) ChannelCat ViT-Small 71.1 34.3
(g) Cross-att DEBiT-b 0.0 0.0
(h) [8] Cross-att DEBiT-b 90.5 60.3
Table 1: Comparisons of agents with different binocular visual encoders, trained and validated with Sliding=True. s2d=SpaceToDepth, not applicable to DEBiT and ViT. PT=pre-training for relative pose and visibility estimation.
@@
Metrics – Navigation performance is evaluated by success rate (SR), i.e., fraction of episodes terminated within a distance of ${<}1$m to the goal by the agent calling the STOP action, and SPL [2], i.e., SR weighted by the optimality of the path, $\textit{SPL}=\frac{1}{N}\sum_{i=1}^{N}S_{i}\frac{\ell_{i}^{*}}{\max(\ell_{i},\ell_{i}^{*})},$ where $S_{i}$ be a binary success indicator in episode $i$, $\ell_{i}$ is the agent path length and $\ell_{i}^{*}$ the shortest path length.
@@
Comparing the comparable – An important choice in the Habitat simulator is the binary Sliding switch, which, when switched on, allows the agent in simulation to slide along obstacles when colliding, against the more realistic behavior of stopping when colliding. This parameter is known to have a big impact on sim2real transfer [33]. In our own analysis, we saw that the increased difficulty of Sliding=False goes beyond sim2real transfer and also impacts task difficulty in simulation: when the agent hits a wall, it can in principle continue the episode, but is frequently restricted in its motion, often stuck, and typically needs to turn to recover.
@@
While this setting is switched to True as a default, there is consensus in the field that it should be set to False, as the goal is to decrease the sim2real gap and to evaluate agents in situations which are as close as possible to real physical robots and environments. Looking closer into the available code of ImageNav repositories, it appears most use Sliding=False (DEBiT [8], ZSEL [1], ZSON [44], PIRLNav [53]), with a couple of exceptions using Sliding=True, like FGPrompt [67]. We therefore performed experiments with both settings, and were surprised to see that there is a big influence of this parameter. Given its impact, we color-coded the entries in all tables into three different groups: (i) Sliding=True: the visual encoder $\phi$ has been trained with this setting on; (ii) Sliding=False: it has been trained with this setting off; (iii) Pre-train: the visual encoder $\phi$ has been pre-trained on an auxiliary task, typically relative pose and visibility estimation (RPVE).
Model s2d Backbone PT SR(%) SPL(%)
(a) Late Fusion ResNet9 12.8 7.1
(b) [1] Late Fusion ResNet9 13.2 8.9
(c) Late Fusion ViT-Small 6.9 4.5
(d) ChannelCat ResNet9 44.6 23.4
(e) [67] ChannelCat ResNet9 31.7 18.7
(f) ChannelCat ViT-Small 35.3 16.2
(g) Cross-att DEBiT-b 0.0 0.0
(h) [8] Cross-att DEBiT-b 81.7§ 52.0§
[…]

### v3

@@ Abstract
Image goal navigation requires two different skills: firstly, core navigation skills, including the detection of free space and obstacles, and taking decisions based on an internal representation; and secondly, computing directional information by comparing visual observations to the goal image. Current state-of-the-art methods either rely on dedicated image-matching, or pre-training of computer vision modules on relative pose estimation. In this paper, we study whether this task can be efficiently solved with end-to-end training of full agents with RL, as has been claimed by recent work. A positive answer would have impact beyond Embodied AI and allow training of relative pose estimation from reward for navigation alone. In this large experimental study we investigate the effect of architectural choices like late fusion, channel stacking, space-to-depth projections and cross-attention, and their role in the emergence of relative pose estimators from navigation training. We show that the success of recent methods is influenced up to a certain extent by simulator settings, leading to shortcuts in simulation. However, we also show that these capabilities can be transferred to more realistic setting, up to some extent. We also find evidence for correlations between navigation performance and probed (emerging) relative pose estimation performance, an important sub skill.
@@ 1 Introduction
Early works addressed the problem in a purely data-driven way by end-to-end training Late Fusion (eg. Siamese) image encoders [80], with mixed results. It was conjectured, that the weak and sparse learning signal from RL might not be sufficient to train high capacity encoders comparing images. Since the ImageNav task introduces a sub task of relative pose estimation (“where is the goal with respect to my current position?”), there is a strong case for modular methods separating perception and decision taking. For instance, Krantz et al. [39] employ local features and matching (SuperGlue [62]) combined with map and plan strategies for navigation. However, it has since been shown [8] that image correspondences can actually emerge directly in the cross-attention layers of a binocular transformer, provided it is appropriately pre-trained. This strategy, used in the DEBiT agent [8] outperforms explicit matching approaches. Key to this solution is, firstly, pre-training, and secondly, the fusion of observation and goal such that comparisons between local parts (e.g. patches) of the two respective images can be calculated by the architecture, as opposed to comparing embedding vectors.
Recent work [70] however indicates that ImageNav can be addressed by very low-capacity convolutional encoders with channel-wise stacking and through RL training alone, without any pre-training. FGPrompt [70] achieves results which are on par with DEBiT with a fraction of the model capacity, which raises interesting questions:
Q1:
Can ImageNav be learned by RL alone?
Q2:
Which architecture can support that best?
Q3:
Is the solution related to the estimation of directional information between observation and goal images?
The implications could potentially go beyond image goal navigation and have impact on the wider field of relative pose estimation and localization. In this paper we systematically disentangle the different mediators to shed a light on what choices matter in ImageNav, and we find that:
R1:
Agents trained with RL alone underperform in realistic navigation settings.
R2:
Early Fusion performs better, with early patch-wise fusion being essential, compared to Late Fusion.
R3:
We find correlations between navigation and (emergent) relative pose estimation performance, which we probe from representations trained with RL.
Bonus:
We discovered that success of recent frugal architectures using channel stacking and trained with RL alone is mostly due to a simulator setting that allows agents to slide along walls. Surprisingly, while this setting is known to hamper sim2real transfer since [33], we have obtained further insights: (i) unrealistic motion simulation has also a negative impact on the learned perception capabilities, and (ii) some learned capabilities can still be transferred to realistic settings if the transfer includes weights of the perception module.
@@ 2 Related Work
Navigation from visual observations – has been addressed in several different fields. Robotics for a long time focused on explicit modeling [13, 45, 47], which is mainly based on mapping and localization [12, 41, 71], explicit planning [38, 66] and low-level control [26, 61]. Accurate sensor and observation models are essential, followed by filtering, dynamical models and optimization techniques.
Modern AI-based solutions are typically trained on large-scale photorealistic simulators like Habitat [63] or AI2-Thor [37]. Modular agents [17, 54, 57] decompose the problem in sub modules, typically mapping, pose estimation, planning and local decision taking. On the other end of this spectrum, end-to-end trained models directly map input to actions, with RL [31, 50, 79, 72] or Imitation Learning (IL) [23], offline-RL [69], or unsupervised RL [36, 44]. In recent work, end-to-end trained agents are combined with visual foundation models, eg. DINOv2 [52] in PoliFormer [79], or binocular ones for image comparisons as in DEBiT [8]. Agent memory is implemented as recurrent representations [10, 32], occupancy maps [17], semantic maps [16], latent metric maps [5, 30, 53], topological maps [6, 18, 67], scene graphs [68], explicit episodic memory [20, 24, 25, 58], implicit representations [48, 40] or navigability [9]. In this work we investigate end-to-end trained agents with recurrent memory.
@@
Relative pose estimation (RPE) – was tackled for decades with pixel-level image matching techniques [28, 51, 64]. Learning-based approaches have also been proposed [34], and to overcome the lack of generalizability of early methods, self-supervised techniques have been used [43]. More recently, DUSt3R [73] regresses pointmaps of each image expressed in the coordinate system of the first image while MASt3R [42] additionally learns an extra descriptor inspired by standard image matching. Both leverage CroCo pre-training [74, 75]. Similarly, MicKey [4] regresses pointmaps of each image in its respective coordinate system and learns descriptors to obtain correspondences, supervised by relative pose  [7, 11]. These recent methods have led to impressive results, even under scenarios with little image overlap as in the MapFree-Relocalization benchmark [3]. In this work, we study RPE as a sub-task of ImageNav and probe it explicitly from agent representations.
@@ 3 Can we learn goal-perception with actions?
We study the ImageNav task in photo-realistic 3D environments, where an agent is given a goal image $\mathbf{g}{\in}\mathbb{R}^{3{\times}H{\times}W}$ and is required to navigate from a starting location to the position shown in the goal. At each timestep $t$ the agent observes an RGB image $\mathbf{o}_{t}{\in}\mathbb{R}^{3{\times}H{\times}W}$, all images are of size $112{\times}112$. The action space is the discrete set $\mathcal{A}$ ={MOVE FORWARD 0.25m, TURN LEFT $10^{\circ}$, TURN RIGHT $10^{\circ}$, and STOP}. An episode is considered successful if the agent is within 1m of the goal position and it calls the STOP action within its 1000 steps budget. We use the Habitat simulator [63] and the Gibson dataset [77].
All our experiments are done with different variants of the same identical base agent, which maintains a recurrent episodic memory $\mathbf{h}_{t}$, integrates the observation $\mathbf{o}_{t}$ and the goal image $\mathbf{g}$, and predicts actions $\mathbf{a}_{t}$:
@@
where $h$ is the function updating the hidden state $\mathbf{h}_{t}$ of a GRU [21] and gating equations have been omitted for brevity; $\phi$ and $\zeta$ are trainable encoders, and $\pi$ is a linear policy. All agent variants are trained from scratch with PPO [65] with reward function defined as in [19, 8]: $r_{t}=\mathrm{K}\cdot\mathbf{1}_{\text{success}}-\Delta_{t}^{\mathrm{Geo}}-\lambda,$ where $\lambda{=}0.01$ is a slack cost to encourage efficiency, K=10K{=}10, and $\Delta_{t}^{\mathrm{Geo}}$ is the increase in geodesic distance to the goal.
@@ 3.1 Architecture design choices
At the heart of the problem is the sub task of learning to compare images $\mathbf{o}_{t}$ and $\mathbf{g}$ to infer directional information related to relative pose estimation. In the classical ImageNav task, goal images can be taken from any position and do not necessarily show a semantically meaningful object. Given the possible absence of semantic cues, which would have allowed to compare images on a global level, we conjecture that successful comparison requires the computation of some form of correspondence between local parts of the respective images. While pose estimation methods based on local feature matching perform this explicitly [59, 62], recent work have shown that this kind of correspondence can emerge implicitly from large-scale training without explicit supervision of matches, e.g. in CroCo [74] or DUSt3R [73].
@@
Assuming that correspondence computations are necessary to extract directional information, we conjecture that these computations need to be supported by underlying network architectures. The basic building blocks for image comparison in the literature, also illustrated in Fig. 2, are:
@@
use separate networks $\phi_{o}$ and $\phi_{g}$ to encode observation and goal. The representation fed to the agent is given as $\phi(\mathbf{o}_{t},\mathbf{g})=[\phi_{o}(\mathbf{o}_{t}),\phi_{g}(\mathbf{g})]$, where $[.]$ denotes concatenation, see Fig. 2(a). The comparison is thus performed “late” between embeddings $\phi_{o}(\mathbf{o}_{t})$ and $\phi_{g}(\mathbf{g})$, which makes it generally harder to be done on a local image level, unless output representations retain sufficient spatial structure going through encoder layers.
@@
uses a single network to encode both observation and goal, which are channel stacked into one input image $\phi\left([\mathbf{o}_{t},\mathbf{g}]_{\textrm{dim}=1}\right)$, where 0 is the batch dimension and 1 the channel dimension — see Fig. 2(b).
If encoders are implemented as CNNs, a common choice in robotics for computational complexity reasons, comparisons between local parts of images can be done only if the receptive field of the network is big enough to encompass the full image. Given a sufficiently deep encoder correspondence can therefore be computed, but not necessarily efficiently, as individual matching results need to be accumulated in the output of different filter kernels.
If encoders are ViTs, each feature patch corresponds to a pair of patches having the same coordinates in both images. Correspondence can be more easily computed if different self-attention heads are learned with query and key projections focusing on one of the two channel groups of the two respective images. [^1] As an illustrative example, consider an input image $X$ corresponding to channel stacking of two images $X_{A}$ and $X_{B}$. Then, if for a given attention head the Query projection only uses the channels of image $X_{A}$ and the Key projection only use the channels of image $X_{B}$, then this self-attention on $X$ mimics cross-attention between $X_{A}$ and $X_{B}$.
@@
reshapes image patches into channel values [60]. FGPrompt [70] uses SpaceToDepth in combination with ChannelCat of observation and goal, achieving strong navigation results. We investigate whether a ResNet in this configuration can compute correspondences across a large spatial dimension through a single convolutional layer, somewhat reminiscent of cross-attention, only with very few parameters — see Fig. 2(c).
@@
is a natural way to compute correspondences between local parts of images [8], as each patch in one image can be naturally linked to one or more patches in the other image through the cross-attention distribution, see Fig. 2(d).
@@ 3.2 Implementations in prior work
Early methods used Late Fusion approaches based on CNNs [1, 80], but later works switched to ViTs. Cross-attention for ImageNav was introduced in DEBiT [8] with code available, [^2] https://github.com/naver/debit and combined with pre-training, first for cross-view completion as in [74], then for relative pose and visibility estimation. The binocular ViT was combined with an additional CNN with only the observation as input, trained with RL, providing current state-of-the-art results. Interestingly, the authors have shown that pre-training is essential, as the high-capacity binocular transformer cannot be trained from scratch with RL only. The learning signal from the RL loss alone seems to be too weak to drive the encoder to discover correspondence computations. The inability of RL to train big encoders was somewhat confirmed in the Late Fusion method OVRL-v2 [78], as “switching from the 50.9M parameters ViT-Small to the 179.2M parameters ViT-Base produces another negative result: SR drops -2.4% while SPL minimally increases by +0.7%”.
FGPrompt [70] proposed channel stacking for ImageNav, with code available, [^3] https://github.com/XinyuSun/FGPrompt and it is based on previous code of ZSEL [1]. Both codebases use a non-standard ResNet9 architecture with several custom blocks, the most significant one being SpaceToDepth described earlier. Interestingly, SpaceToDepth was first introduced in TResNet [60] as a general purpose network block for computer vision optimized for memory efficiency, targeting image classification and object detection. Correspondence problems were not explicitly targeted. The ResNet9 encoder in this paper is exactly the same as those used in ZSEL and FGPrompt.
Given the good results FGPrompt obtained without any pre-training, 92% SR, vs. DEBiT’s 94% SR relying on pre-training, we raise the question whether it is possible to use the weak learning signal provided by RL to learn binocular perception modules capable of providing directional information. A positive answer would have significant impact also on training vision modules targeting RPE directly. In the experimental section we will dive into this question and benchmark architectures and design choices.
@@ 4 Experimental Setup
Unfortunately there are different evaluation protocols in the literature on Embodied AI, and in particular on the ImageNav task. We therefore train all agents ourselves in identical conditions and with the same experimental protocol. The agents are trained on the 72 scenes of the Gibson dataset [77] and use the standard Habitat episode definitions (a subset of the literature uses the definitions from Episodic Transformers [49] and another subset those from NSNRL [27]). We set the maximum episode length to the default value of 1000 steps and we train for 500M steps, which allows all agent variants to converge easily.
@@
Implementation details – We test perception networks $\phi$ implemented as ResNet9, ViT-Small [^4] https://github.com/huggingface/pytorch-image-models and DEBiT-B architectures, available in their respective public repositories. We test CannelCat, Late Fusion and SpaceToDepth variants of ResNet and ViT, while for DEBiT we either use the pretrained binocular encoder or train the whole agent from scratch. Network versions are chosen based on size-performance tradeoff — see Supplementary for details. The function $\zeta$ embeds previous actions into a 32D feature, $h$ is a GRU with 2 layers and hidden dimension 128, followed by a linear Actor-Critic policy $\pi$.
@@
Validation – To avoid overfitting evaluation over the choice of checkpoints, while at the same time staying comparable, we use three different splits. We train on the standard Gibson train split (72 scenes) and report on the standard Gibson validation split (14 scenes), as most literature, but we use an independent custom holdout set of unused Gibson scenes to choose the checkpoint.
@@
Metrics – Navigation performance is evaluated by success rate (SR), i.e., fraction of episodes terminated within a distance of <1{<}1m to the goal by the agent calling the STOP action, and SPL [2], i.e., SR weighted by the optimality of the path, $\textit{SPL}=\frac{1}{N}\sum_{i=1}^{N}S_{i}\frac{\ell_{i}^{*}}{\max(\ell_{i},\ell_{i}^{*})},$ where $S_{i}$ be a binary success indicator in episode $i$, $\ell_{i}$ is the agent path length and $\ell_{i}^{*}$ the shortest path length.
@@
Comparing the comparable – An important choice in the Habitat simulator is the binary Sliding switch, which, when set to True, allows the agent in simulation to slide along obstacles when colliding, against the more realistic behavior of stopping. This parameter is known to have a big impact on sim2real transfer [33]. In our own analysis, we saw that the increased difficulty of Sliding=False goes beyond sim2real transfer and also impacts task difficulty in simulation: when the agent hits a wall, it can in principle continue the episode, but is frequently restricted in its motion, often stuck, and typically needs to turn to recover.
@@
While this setting is switched to True by default, there is consensus in the field that it should be set to False, as the goal is to decrease the sim2real gap and to evaluate agents in situations which are as close as possible to real physical robots and environments. Looking closer into the available code of navigation repositories, it seems that indeed most use Sliding=False (DEBiT [8], ZSEL [1], ZSON [46], PIRLNav [56]), with a couple of exceptions using Sliding=True, notably FGPrompt [70] (details in Supplementary). We therefore performed experiments with both settings, and were surprised to see that there is a big influence of this parameter. Given its impact, we color-coded the entries in all tables into three different groups: (i) Sliding=True: the visual encoder $\phi$ has been trained with this setting on; (ii) Sliding=False: it has been trained with this setting off; (iii) Pre-train: the visual encoder $\phi$ has been pre-trained on an auxiliary task, typically relative pose and visibility estimation (RPVE).
Slide=True Slide=False
Model s2d Backbone SR SPL SR SPL
(a) Late Fusion ResNet9 13.8 8.0 12.8 7.1
(b) Late Fusion [1] ResNet9 12.5 7.6 13.2 8.9
(c) Late Fusion ViT-Small 12.5 6.7 6.9 4.5
(d) ChannelCat ResNet9 83.2 43.9 44.6 23.4
(e) ChannelCat [70] ResNet9 83.6 42.1 31.7 18.7
(f) ChannelCat ViT-Small 71.1 34.3 35.3 16.2
(g) Cross-attn DEBiT-B 0.0 0.0 0.0 0.0
(h) Cross-attn [8] DEBiT-B 90.5 60.3 81.7 52.0
Table 1: Agents with different visual encoders trained and validated with Sliding=True or Sliding=False. s2d=SpaceToDepth. $\ddagger$=pre-trained for RPVE.